If I roll a standard 6-faced die three times, what is the probability that I roll at least one 6? Express your answer as a common fraction.

Answer by stanbon(75887)

If I roll a standard 6-faced die three times, what is the probability that I roll at least one 6? Express your answer as a common fraction.
----
General rule: P(none) + P(at least one) = 1
----
Your Problem:
P(no 6 in three rolls) = (5/6)^3
-----
P(at least one 6 in three rolls) = 1 - (5/6)^3 = 0.4213..
Cheers,
Stan H.